Circular No. 8397

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ed IN NGC 6786 M. Armstrong (cf. IAUC 8376) and T. Boles (cf. IAUC 8396) inde-

pendently report their discovery of an apparent supernova on unfiltered CCD images. SN 2004ed is located at α = 19^{h}10^{m}53^{s}.62, δ = +73^{o}24′27′′.6 (equinox 2000.0), which is ≈ 1′′.8 west and 9′′.8 south of the center of NGC 6786, according to Armstrong, who adds that nothing is visible at this position on Palomar Sky Survey images from 1993 (limiting red mag 20.8) and 1992 (limiting blue mag 22.5). Boles provides position end figures 53^{s}.72, 27′′.0. Available magnitudes: 2002 July 21, [19.5 (Armstrong); 2004 June 17, [19.5 (Armstrong); July 16, [19.5 (Boles); Sept. 2.980 UT, 17.4 (Armstrong); 3.849, 17.4 (Armstrong); 3.886, 17.5 (Boles).

2002 CE_{26}

M. K. Shepard and J. Schlieder, Bloomsburg University of Pennsylvania; M. C. Nolan and A. A. Hine, National Astronomy and Ionosphere Center; and L. A. M. Benner, S. J. Ostro, and J. D. Giorgini, Jet Propulsion Laboratory, report: “Arecibo delay-Doppler radar images (2380 MHz, 12.6 cm) obtained on Aug. 30–31 show that 2002 CE_{26} is a binary system. Preliminary estimates of average diameters, based on range estimates at 75-m resolution, are 3 and 0.2 km. The orbit was seen at approximately superior conjunction at Aug. 30.076 UT and inferior conjunction at Aug. 31.052. The average orbital radius, based on these and Doppler-bandwidth observations, is 5 km. Preliminary estimates, based on apparent motion of the satellite in range, indicate an orbital period of ≈ 16 hours.”

SUPERNOVA 2004dj IN NGC 2403 P. Chandra and A. Ray, Tata Institute of Fundamental Research,

Mumbai, report the detection of radio emission from SN 2004dj with the Giant Metrewave Radio Telescope (GMRT) in the L band on Aug. 12 and 22. The flux density on Aug. 12 at the center frequency of 1390 MHz was 300 ± 85 μJy, which grew to 900 ± 105 μJy by Aug. 22. On Aug. 22, SN 2004dj remained undetected in the GMRT 1060-MHz band up to a 3σ limit of 350 μJy. Total bandwidths of 32 MHz were used at both frequencies for each observation. GMRT observations in the L band are continuing.

CCD R magnitudes by K. Hornoch, Lelekovice, Czech Rep.: Aug. 4.047, 11.55 (poor sky conditions); 10.072, 11.60; 16.076, 11.60; 19.042, 11.59; 23.029, 11.61; 28.985, 11.64.

2004 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green